I Need Help please :-(!!!!
What is the probability that at least 3 red cards will be drawn from a 52 card deck?

Answer by jrfrunner(365)

It depends on a couple of things
1. How many cards are you selecting, for which you want to know P(at least 3 red)
2.with replacement or without replacement
You are asking about selecting 4 cards without replacement and probability of selecting X=at least 3 red cards
There are 26 red cards in a deck of 52 cards
choosing 3 red out of 4 cards =RRRB but there are 4C3 ways to choose that black one
Choosing 4 reds out of 4 cards = RRRR
P(X>=3)=P(X=3)+P(X=4)= 4C3*P(RRRB)+4C4*P(RRRR)=4C3*(26/52)*(25/51)*(24/50)*(26/49) + 4C4*(26/52)*(25/51)*(24/50)*(23/49)
= 0.2497 + 0.0552=0.30492
another way to solve this is
P(X>=3)=P(X=3)+P(X=4)= (26C3*26C1)+26C4*26C0)/52C4